# A Survey on Failure Analysis and Fault Injection in AI Systems

### v1

@@ 3. Survey Methodology
In total, our recursive search methodology enabled us to collect 160 publications and code repositories about FA and FI in AI systems, spanning from the year 2001 to 2024. These publications and repositories are classified into six layers by research topics including AI service, AI model, AI framework, AI toolkit, AI platform, and AI infrastructure. Moreover, the distribution of different classes is presented in Fig.4. It is important to note that some papers may address more than one research topic. Consequently, the total number of papers and repositories in Fig.4 is larger than 160. Next, we will show the details on FA and FI in different layers of AI systems.
Figure 3. Overall life cycle of an AI system failure.
@@
Figure 4. Distribution of failure analysis and fault injection paper and repositories across layers.
@@ 4. Failure Analysis and Fault Injection in AI Service
In the context of AI systems, the AI service layer can be considered as the topmost layer that directly interacts with users or other systems. This layer acts as an interface or entry point for users to access and utilize the AI capabilities provided by the underlying layers of the AI system. Potential failures at this layer can include unavailability of the service or incorrect outputs, such as incorrect inference from a classifier or hallucinations from an LLM-based service. This section delves into the FA and FI in AI services
@@ 4.1. Failure Analysis in AI Service
Data Fault. These faults related to the format, type, and noise of data can lead to the failure of AI services (Schelter et al., 2021; Tan et al., 2023; Zhao et al., 2022). For example, incorrect data encoding (e.g., requesting UTF-8 but receiving ASCII) or inappropriate data types (e.g., expecting a string but receiving an integer) can prevent AI services from working properly. The JENGA framework explores the impact of data faults on predictions of machine learning models (Schelter et al., 2021). Furthermore, data drift and concept drift are common problems (Tan et al., 2023). Data drift refers to a model trained on a specific distribution of data but then encountering a different distribution in practice. While concept drift occurs as the relationship between features and labels becomes invalid over time. Zhao et al. (Zhao et al., 2022) investigate the impact of concept drift on model performance.
@@
Development Fault. The primary reason for service failures is code quality such as bugs, logical faults in code, and poor software design (Lenarduzzi et al., 2021; Gezici and Tarhan, 2022; Hu et al., 2021). Code faults typically originate from coding mistakes. Logical faults in code often involve incorrect algorithm implementation, impacting the accuracy and efficiency of inference. Thus, poor software design affects system performance. Additionally, failures in AI service API calls (Wan et al., 2021), originating from API incompatibility, API change, and API misuse, can lead to AI service failures.
@@
Deployment Fault. During the deployment of AI services, various faults can arise, including outdated models (Mousavi et al., 2024), path configuration errors (Hien Luu, 2024), and inappropriate resource allocation (Mao et al., 2023). These faults can impact system performance and stability. As data changes over time, model performance may degrade, necessitating regular updates and retraining to maintain their effectiveness. Path configuration faults can prevent the proper loading of models and data. Inadequate resource allocation (Mao et al., 2023), especially inefficient use of CPU and GPU resources, can lead to decreased system performance and unnecessary waste.
@@
Network Transmission Fault. It may arise from network congestion, bandwidth limitations, or failures in network hardware, leading to packet delays or losses (Lamberti, 2023). Network congestion occurs when data traffic exceeds the network’s bandwidth capacity, preventing data transmission on time. Bandwidth limitations are imposed by the maximum transmission rate of a network connection, often determined by service provider or the capabilities of network hardware. Furthermore, physical damage or configuration faults in network devices can also lead to packet delays or losses. failures in network hardware, leading to packet delays or losses (Lamberti, 2023). Network congestion occurs when data traffic exceeds the network’s bandwidth capacity, preventing data transmission on time. Bandwidth limitations are imposed by the maximum transmission rate of a network connection, often determined by service provider or the capabilities of network hardware. Furthermore, physical damage or configuration faults in network devices can also lead to packet delays or losses.
@@
External Attack. In addition to internal faults within AI services, external attacks can also lead to service failures. These include network attacks such as Distributed Denial of Service (DDoS) attacks (Matti and Fredrik, 2023) and Man-in-the-Middle (MITM) attacks (Rao et al., 2023), which can not only cause temporary service interruptions but also lead to data leakage or corruption. Moreover, adversarial attacks target the AI model by designing malicious inputs (e.g., meticulously modified images or texts) to deceive the AI into making incorrect actions (Abdelnabi et al., 2023; Akhtar et al., 2021; Qi et al., 2024). Adversarial attacks further divide into white-box attacks and black-box attacks. In white-box attacks, the deployed model is fully understood, including inputs and architecture, allowing for targeted attacks. In black-box attacks, only the model’s inputs and output labels/confidences are accessible.
@@
Development Defective Code Logical faults in code, and poor software design. (Lenarduzzi et al., 2021; Gezici and Tarhan, 2022; Hu et al., 2021)
@@
External Attack Adversarial Attack Use known information and relevant patterns to attack the model. FGSM (Goodfellow et al., 2015), PGD (Madry et al., 2018), DI-AA (Wang et al., 2022), Grey-box attack (Lapid and Sipper, 2023)
@@ 4.2. Fault Injection in AI Service
Fault injection in AI service layer encompasses four dimensions including data, service API, network transmission, and external attack. By simulating diverse fault scenarios in these dimensions, it is possible to assess the system’s robustness and reliability, thereby preventing inaccurate predictions or even system failures.  Table 2 illustrates the current FI tools in the AI service layer, followed by a description of each of them.
Data Fault. Data perturbation is the most intuitive method of FI at the data dimension. Introducing noise or modifying data manually can simulate uncertainties in real-world data. It can be achieved using tools such as NumPy (Numpy, 2024), Scikit-learn (Scikit-learn, 2024), and so on. JENGA (Schelter et al., 2021) is a framework that studies the impact of data faults (e.g., missing values, outliers, typing errors, and noisy) on the predictions of machine learning models. Additionally, data and concept drift can be simulated through carefully designed data disturbances. Moreover, tools such as scikit-multiflow (Montiel et al., 2018) and MOA (Bifet et al., 2010) enable the simulation of sudden, gradual, and incremental drifts in data streams.
@@
External Attack. Adversarial attacks are deliberate attacks on the AI service, including both white-box and black-box attacks. White-box attacks leverage model structure and parameter information to strategically generate adversarial samples, such as FGSM (Goodfellow et al., 2015), PGD (Madry et al., 2018), and DI-AA (Wang et al., 2022). Black-box attacks lack insight into the model and rely on observing model inputs and outputs to create adversarial samples, such as HopSkipJump (Chen et al., 2020b), PopSkipJump (Simon-Gabriel et al., 2021), and GeoDA (Rahmati et al., 2020). Grey-box attacks utilize partial information to generate adversarial samples. Raz et al. (Lapid and Sipper, 2023) attack image-to-text models based on adversarial perturbations. In addition to traditional methods, prompt fault injection has become popular for large models. It guides models to generate special outputs by adding prompts to the input text. Greshake et al. (Abdelnabi et al., 2023) use indirect prompt injection to exploit LLM-integrated applications and systematically investigate impacts and vulnerabilities, including data theft, worming, information ecosystem contamination, and other novel security risks. PromptAid (Mishra et al., 2023) can introduce keyword and paraphrasing perturbations into prompts. HouYi (Liu et al., 2023a) conducts prompt fault injections in a black-box manner. Zhang et al. (Zhang et al., 2024) propose goal-guided generative prompt injection attack on LLMs.
@@ 4.3. Gap between Failure Analysis and Fault Injection in AI Service
Failure JENGA MOA FGSM PromptAid Istio ChaosBlade Covered
Data Quality True
@@
Defective Code False
@@
Configuration Fault False
Network Transmission Fault True
Network Attack True
@@
Incomplete Coverage. The scrutinized FI tools, in aggregate, encompass a substantial proportion of the delineated fault types. Nonetheless, certain fault types (e.g., ”Defective Code”) remain unaddressed by any of the existing tools. This revelation underscores the exigency for subsequent research to ameliorate these deficiencies in fault simulation.
@@
New Emerging Fault Types. The incorporation of fault types such as ”Prompt Injection Attacks” accentuates the burgeoning significance of accounting for external factors and security facets in the FA of AI services. Nevertheless, traditional FI tools (e.g., ChaosBlade) are designed for cloud computing services and do not consider the specific scenarios of AI systems. As AI services become increasingly interconnected and susceptible to a myriad of threats, it is imperative to cultivate FI tools capable of efficaciously simulating failures emanating from these nascent domains.
@@ 5.1. Failure Analysis in AI Model
Dataset Partitioning Fault Insufficient data for training and validation. (Muraina, 2022)
Data Fault. The quality of training data is pivotal for successful model training. Alice et al. (Croft et al., 2023) conduct a systematic analysis of data quality attributes (accuracy, uniqueness, consistency, completeness, and timeliness) across five software bug datasets and find that 20-71% of data labels are inaccurate, which could severely hinder model training in extreme cases. Some studies (Whang et al., 2023; Liang et al., 2022) have examined challenges faced by data quality. Furthermore, data preprocessing and augmentation are crucial. Raw data are susceptible to noise, damage, loss, and inconsistency, thus necessitating preprocessing steps (e.g., data cleaning, integration, transformation, and reduction) to facilitate easier knowledge extraction from datasets. Data augmentation aims to expand the training dataset through specific transformations to enhance the model’s generalizability. Das (Das, 2022) lists ten common faults in data preprocessing, while Maharana et al. (Maharana et al., 2022) discuss various data preprocessing and data augmentation techniques to enhance model performance.
@@
Model Structure and Algorithm Fault. Shiri et al. (Shiri et al., 2023) investigate various aspects of different models and evaluate their performance on three public datasets. Activation functions are crucial for introducing non-linearity. Dubey et al. (Dubey et al., 2022) compare the performance of 18 distinct activation functions (e.g., Sigmoid, Tanh, and ReLU) on various datasets. Model training requires regularization to avoid overfitting. Tian et al. (Tian and Zhang, 2022) compare different regularization techniques, including sparse regularization, low-rank regularization, dropout, batch normalization, and others. They discusse the selection of regularization techniques for specific tasks. The selection of optimizers significantly affects model performance and training speed. Haji et al. (Haji and Abdulazeez, 2021) compare various optimizers like SGD, Adam, AdaGrad, etc. They highlight the advantages and disadvantages of these optimizers in terms of training speed, convergence rate, and performance. The loss function is also essential for minimizing the discrepancy between predicted results and target values. Wang et al. (Wang et al., 2020) introduce 31 loss functions from five aspects: classification, regression, unsupervised learning of traditional machine learning, object detection, and face recognition of deep learning. Additionally, the ratio of training to validation data must be considered to ensure the model has enough data for learning while the validation data is adequate for model adjustments (Muraina, 2022).
@@ 5.2. Fault Injection in AI Model
Recent studies in mutation testing have made significant contributions. We have summarized these works as shown in Table 5. DeepMutation (Ma et al., 2018) designs 13 mutation operators that inject faults into both the training data and code of deep learning models. DeepMutation++ (Hu et al., 2019) combines DeepMutation (eight model-level operators for FNN models) and proposes nine new operators specifically for RNN models, enabling both static mutations in FNNs and RNNs and dynamic mutations in RNNs. MuNN (Shen et al., 2018) develops five mutation operators, focusing on model-level fault injection. DeepCrime (Humbatova et al., 2021) implements 24 deep learning mutation operators to test the robustness of deep learning systems, including training data operators, hyperparameters operators, activation function operators, regularization operators, weights operators, loss function operators, optimization operators, and validation operators. Additionally, some studies have focused on mutation testing in reinforcement learning (Tambon et al., 2023; Lu et al., 2022b) and unsupervised learning (Lu et al., 2022a).

### v2

@@ 3.3. Summary of Our Survey
These studies are classified into six layers by research topics including AI service, AI model, AI framework, AI toolkit, AI platform, and AI infrastructure. Moreover, the distribution of different classes is presented in Fig. 6. It is important to note that some studies may address more than one research topic. Consequently, the total number of studies in Fig. 6 is larger than 142. Next, we will show the details on FA and FI in different layers of AI systems.
Figure 5. Distribution of studies across layers.
@@
Figure 6. Distribution of failure analysis and fault injection studies across layers.
3.4. Comparing Against Existing Surveys
There are three existing works (Natella et al., 2016; Gezici and Tarhan, 2022a; Martínez-Fernández et al., 2022) that are related to our survey.
  • Silverio et al. (Martínez-Fernández et al., 2022) conducted a survey on the literature in the field of Software Engineering for Artificial Intelligence. They found that the most studied characteristics of AI systems are reliability and security. Research related to software testing and software quality is prevalent, but areas such as software maintenance seem to be overlooked, with fewer related papers. The focus of this paper is on failure analysis, which falls under the umbrella of software maintenance, thereby addressing a gap that was not considered in their article. This highlights the importance and relevance of our survey, as it contributes to a more comprehensive understanding of the challenges and solutions in the field of AI systems.
  • Bahar et al. (Gezici and Tarhan, 2022a) aimed to investigate the recent advancements in Software Quality (SQ) based on AI systems and identify quality attributes, application models, challenges, and practices reported in the literature. Their focus was primarily on how to develop a high-quality AI system. However, they did not delve into the analysis and handling of AI system failures. Our study complements the work of Bahar et al. (Gezici and Tarhan, 2022a) by providing a more comprehensive view of software quality in AI systems, extending beyond development to include maintenance and fault tolerance.
  • Natella et al. (Natella et al., 2016) provided a comprehensive overview of the latest advancements in software fault injection to assist researchers and practitioners in selecting the methods most suitable for their reliability assessment goals. However, their focus was primarily on traditional software systems, not AI systems. As shown in Fig. 2, AI systems differ significantly from previous systems. They have unique characteristics and complexities, such as the use of machine learning models, which introduce new types of fault (Sinha and Lee, 2024; Liu et al., 2022). Therefore, there is a need for research specifically targeted at AI systems.
Therefore, we claim that our study is the first to provide a comprehensive review of the scientific literature in the context of FA and FI specifically for AI-based systems. By focusing on the unique challenges and characteristics of AI systems, our objective is to fill a critical gap in the existing literature and provide valuable insights to researchers and practitioners in this rapidly evolving field. Our survey not only synthesizes the current state of the art, but also identifies key areas for future research, thereby paving the way for the development of more reliable and robust AI systems.
@@ 4. Failure Analysis and Fault Injection in AI Service
In the context of AI systems, the AI service layer can be considered as the top layer that interacts directly with users or other systems. This layer acts as an interface or entry point for users to access and utilize the AI capabilities provided by the underlying layers of the AI system. Potential failures at this layer can include unavailability of the service or incorrect output, such as incorrect inference from a classifier or hallucinations from an LLM-based service. This section delves into the FA and FI in AI services
@@ 4.1. Failure Analysis in AI Service
Data Fault. These faults related to the format, type, and noise of the data can lead to the failure of AI services (Schelter et al., 2021; Tan et al., 2023; Zhao et al., 2022). For example, incorrect data encoding (e.g., requesting UTF-8 but receiving ASCII) or inappropriate data types (e.g., expecting a string but receiving an integer) can prevent AI services from working properly. The JENGA framework explores the impact of data faults on predictions of machine learning models (Schelter et al., 2021). Furthermore, data drift and concept drift are common problems (Tan et al., 2023). Data drift refers to a model trained on a specific distribution of data but then encountering a different distribution in practice. However, concept drift occurs as the relationship between features and labels becomes invalid over time. Zhao et al. (Zhao et al., 2022) investigate the impact of concept drift on model performance.
@@
Development Fault. The primary reason for service failures is code quality such as bugs, logical faults in code, and poor software design (Lenarduzzi et al., 2021; Gezici and Tarhan, 2022b; Hu et al., 2021). Code faults typically originate from coding mistakes. Logical faults in code often involve incorrect algorithm implementation, impacting the accuracy and efficiency of inference. Thus, poor software design affects system performance. Additionally, failures in AI service API calls (Wan et al., 2021), originating from API incompatibility, API change, and API misuse, can lead to AI service failures.
@@
Deployment Fault. During the deployment of AI services, various faults can arise, including outdated models (Mousavi et al., 2024), path configuration errors (Hien Luu, 2024), and inappropriate resource allocation (Mao et al., 2023). These faults can impact system performance and stability. As data change over time, model performance may degrade, necessitating regular updates and retraining to maintain their effectiveness. Path configuration faults can prevent the proper loading of models and data. Inadequate resource allocation (Mao et al., 2023), especially inefficient use of CPU and GPU resources, can lead to decreased system performance and unnecessary waste.
@@
Network Transmission Fault. Failures in network hardware leading to packet delays or losses (Lamberti, 2023). Network congestion occurs when data traffic exceeds the network’s bandwidth capacity, preventing data transmission on time. Bandwidth limitations are imposed by the maximum transmission rate of a network connection, often determined by service providers or the capabilities of network hardware. Furthermore, physical damage or configuration faults in network devices can also lead to packet delays or losses.
@@
External Attack. In addition to internal faults within AI services, external attacks can also lead to service failures. These include network attacks such as Distributed Denial of Service (DDoS) attacks (Matti and Fredrik, 2023) and Man-in-the-Middle (MITM) attacks (Rao et al., 2023), which can not only cause temporary service interruptions but also lead to data leakage or corruption. Moreover, adversarial attacks target the AI model by designing malicious inputs (e.g., meticulously modified images or texts) to deceive the AI into making incorrect actions (Abdelnabi et al., 2023; Akhtar et al., 2021; Qi et al., 2024). Adversarial attacks are divided into white-box attacks and black-box attacks. In white-box attacks, the deployed model is fully understood, including inputs and architecture, allowing for targeted attacks. In black-box attacks, only the model’s inputs and output labels/confidences are accessible.
@@
Development Defective Code Logical faults in code, and poor software design. (Lenarduzzi et al., 2021; Gezici and Tarhan, 2022b; Hu et al., 2021)
@@
External Attack Adversarial Attack Use known information and relevant patterns to attack the model. FGSM (Goodfellow et al., 2015), PGD (Madry et al., 2018), DI-AA (Wang et al., 2022a), Grey-box attack (Lapid and Sipper, 2023)
@@ 4.2. Fault Injection in AI Service
Fault injection in AI service layer encompasses four dimensions, including data, API service, network transmission, and external attack. By simulating diverse fault scenarios in these dimensions, it is possible to assess the system’s robustness and reliability, thereby preventing inaccurate predictions or even system failures.  Table 3 illustrates the current FI tools in the AI service layer, followed by a description of each of them.
Data Fault. Data perturbation is the most intuitive method of FI at the data dimension. Introducing noise or modifying the data manually can simulate uncertainties in real-world data. It can be achieved using tools such as NumPy (Numpy, 2024), Scikit-learn (Scikit-learn, 2024), and so on. JENGA (Schelter et al., 2021) is a framework that studies the impact of data faults (e.g., missing values, outliers, typing errors, and noisy) on the predictions of machine learning models. Furthermore, data and concept drift can be simulated through carefully designed data disturbances. Moreover, tools such as scikit-multiflow (Montiel et al., 2018) and MOA (Bifet et al., 2010) enable the simulation of sudden, gradual, and incremental drifts in data streams.
@@
External Attack. Adversarial attacks are deliberate attacks on the AI service, including both white-box and black-box attacks. White-box attacks leverage model structure and parameter information to strategically generate adversarial samples, such as FGSM (Goodfellow et al., 2015), PGD (Madry et al., 2018), and DI-AA (Wang et al., 2022a). Black-box attacks lack insight into the model and rely on observing model inputs and outputs to create adversarial samples, such as HopSkipJump (Chen et al., 2020b), PopSkipJump (Simon-Gabriel et al., 2021), and GeoDA (Rahmati et al., 2020). Grey-box attacks utilize partial information to generate adversarial samples. Raz et al. (Lapid and Sipper, 2023) attack image-to-text models based on adversarial perturbations. In addition to traditional methods, prompt fault injection has become popular for large models. It guides models to generate special outputs by adding prompts to the input text. Greshake et al. (Abdelnabi et al., 2023) use indirect prompt injection to exploit LLM-integrated applications and systematically investigate impacts and vulnerabilities, including data theft, worming, information ecosystem contamination, and other novel security risks. PromptAid (Mishra et al., 2023) can introduce keyword and paraphrasing perturbations into prompts. HouYi (Liu et al., 2023a) conducts prompt fault injections in a black-box manner. Zhang et al. (Zhang et al., 2024) propose a goal-guided generative prompt injection attack on LLMs.
@@ 4.3. Gap between Failure Analysis and Fault Injection in AI Service
Group Failure JENGA MOA FGSM PromptAid Istio ChaosBlade Covered
Data Data Quality True
@@
Development Defective Code False
Configuration Fault False
@@
Network Network Transmission Fault True
Attack Network Attack True
@@
Incomplete Coverage. The scrutinized FI tools, in aggregate, encompass a substantial proportion of the delineated fault types. Nonetheless, certain fault types (e.g., “Defective Code” ) remain unaddressed by any of the existing tools. This revelation underscores the exigency for subsequent research to ameliorate these deficiencies in fault simulation.
@@
New Emerging Fault Types. The incorporation of fault types such as “Prompt Injection Attacks” accentuates the burgeoning significance of accounting for external factors and security facets in the FA of AI services. Nevertheless, traditional FI tools (e.g., ChaosBlade) are designed for cloud computing services and do not consider the specific scenarios of AI systems. As AI services become increasingly interconnected and susceptible to a myriad of threats, it is imperative to cultivate FI tools capable of effectively simulating failures emanating from these nascent domains.
@@ 5.1. Failure Analysis in AI Model
Dataset Partitioning Fault Insufficient data for training and validation. (Joseph, 2022)
Data Fault. The quality of training data is pivotal for successful model training (Yin et al., 2023). Alice et al. (Croft et al., 2023) conduct a systematic analysis of data quality attributes (accuracy, uniqueness, consistency, completeness, and timeliness) across five software bug datasets and find that 20-71% of data labels are inaccurate, which could severely hinder model training in extreme cases. Some studies (Whang et al., 2023; Liang et al., 2022) have examined challenges faced by data quality. Furthermore, data preprocessing and augmentation are crucial. Raw data are susceptible to noise, damage, loss, and inconsistency, thus necessitating preprocessing steps (e.g., data cleaning, integration, transformation, and reduction) to facilitate easier knowledge extraction from datasets. Data augmentation aims to expand the training dataset through specific transformations to enhance the model’s generalizability. Das (Das, 2022) lists ten common faults in data preprocessing, while Maharana et al. (Maharana et al., 2022) discuss various data preprocessing and data augmentation techniques to enhance model performance.
@@
Model Structure and Algorithm Fault. Shiri et al. (Shiri et al., 2023) investigate various aspects of different models and evaluate their performance on three public datasets. Activation functions are crucial for introducing non-linearity. Dubey et al. (Dubey et al., 2022) compare the performance of 18 distinct activation functions (e.g., Sigmoid, Tanh, and ReLU) on various datasets. Model training requires regularization to avoid overfitting. Tian et al. (Tian and Zhang, 2022) compare different regularization techniques, including sparse regularization, low-rank regularization, dropout, batch normalization, and others. They discuss the selection of regularization techniques for specific tasks. The selection of optimizers significantly affects model performance and training speed. Haji et al. (Haji and Abdulazeez, 2021) compare various optimizers such as SGD, Adam, AdaGrad, etc. They highlight the advantages and disadvantages of these optimizers in terms of training speed, convergence rate, and performance. The loss function is also essential for minimizing the discrepancy between predicted results and target values. Wang et al. (Wang et al., 2020) introduce 31 loss functions from five aspects: classification, regression, unsupervised learning of traditional machine learning, object detection and face recognition of deep learning. Furthermore, the ratio of training to validation data must be considered to ensure that the model has enough data for learning while the validation data is adequate for model adjustments (Joseph, 2022).
@@ 5.2. Fault Injection in AI Model
Recent studies in mutation testing have made significant contributions. We have summarized these works as shown in Table 6. DeepMutation (Ma et al., 2018) designs 13 mutation operators that inject faults into both the training data and code of deep learning models. DeepMutation++ (Hu et al., 2019) combines DeepMutation (eight model-level operators for FNN models) and proposes nine new operators specifically for RNN models, enabling both static mutations in FNNs and RNNs and dynamic mutations in RNNs. MuNN (Shen et al., 2018) develops five mutation operators, focusing on model-level fault injection. DeepCrime (Humbatova et al., 2021) implements 24 deep learning mutation operators to test the robustness of deep learning systems, including training data operators, hyperparameters operators, activation function operators, regularization operators, weight operators, loss function operators, optimization operators, and validation operators. Additionally, some studies have focused on mutation testing in reinforcement learning (Tambon et al., 2023; Lu et al., 2022b) and unsupervised learning (Lu et al., 2022a).